If a sloth moves at 2.6 feet per .8 minutes how many miles per hour does it move?
Answer by Maths68(1474)

Change feet into miles and minutes into hour
5280 feet=1 mile
5280/5280 =1/5280
1=0.000198
2.6*1=2.6 * 0.000198
2.6 feet=0.00051 miles
60 minutes = 1 hour
60/60=1/60 hour
1 = 0.01666
0.8*1=0.8*0.01666
0.8 minutes=0.01333 hours

Rephrase your question:-
If a sloth moves at 0.00051 miles per 0.01333 hours how many miles per hour does it move?
Speed = Distance / time
Speed = 0.00051 / 0.01333
Speed = 0.03825 mph